# BarcodeBERT: Transformers for Biodiversity Analyses

### v2

@@ 4.2.2 Zero-shot image classification using DNA embeddings
Table 2: Evaluation of DNA barcode models in a Bayesian zero-shot learning task on the INSECT dataset. The pretraining and fine-tuning data source is indicated by the respective DNA type, and ‘–’ signifies the absence of training for that type. We also indicate the most specific taxon subset. For the baseline CNN encoder, we report the original paper result (left) and reproduced result (right). Numbers in boldface indicate the best result across each task, and underlined second best.
@@ 4.3 Ablation Studies
We study the impact of the different components involved during pretraining. We use the terminology context tokens for tokens that are left unchanged to provide context to the model during pretraining and the terminology substitution tokens for tokens that will be changed for the masked language modelling task. We consider different strategies to calculate the loss of each group separately. The loss associated with predicting contextual tokens is referred to as the “context component” of the loss [^1] Although predictions in the foundation model literature are typically restricted to substitution tokens, we extend this to include context token predictions, maintaining the terminology to explore their potential utility in input reconstruction similar to non-masked autoencoding objectives., while the loss related to predicting substitution tokens is the “substitution component”. By assigning different weights to these two loss components, we sought to observe how these adjustments would affect both training and evaluation. In particular, we define $w_{s}$ as the penalty weight given to the substitution component of the loss. Note that $1-w_{s}$ is always the weight of the context component of the loss.
@@ 4.3.2 Weight of the substitution component of the loss
Table 3: Classification accuracy over the different substitution token ratios $r_{s}$, while keeping constant all of the model architecture (4-4), the value of $k=4$ during tokenization and the penalty weight for the substitution component of the loss ($w_{s}=1$). Numbers in boldface indicate the highest accuracies and the $\rightarrow$ highlighted $\leftarrow$ value shows the selected optimal parameter.
@@
Finetuned Linear probe 1-NN probe
@@
Table 4: Genus-level accuracy for 1-NN probing of unseen species with varying penalty weight assigned to the substitution component of the loss ($w_{s}$). The model architecture remains fixed (4-4), with substitution token ratio $r_{s}$ = 50%. Two tokenizers were tested: a $k$-mer tokenizer with $k$-mer sizes of 2, 4, 6, and 8, and a BPE tokenizer used in DNABERT-2 with a fixed vocabulary size of 4096. Note that the weight of the context component of the loss always equals $1-w_{s}$. The number in boldface indicates the overall best accuracy, underlined the best per tokenizer, and $\rightarrow$ highlighted $\leftarrow$ the selected optimal parameter.
@@ 4.3.3 Tokenization strategies
We evaluated the BPE tokenizer, which generates variable-length tokens based on character co-occurrence frequencies (Sennrich et al., 2016). BPE was designed for subword tokenization to overcome fixed vocabulary (Sennrich et al., 2016) by compressing sequences to reduce the information required for representation (Gallé, 2019). Unlike overlapping $k$-mers, BPE has an advantage for masked DNA language models (Zhou et al., 2023), since it avoids information leaks from adjacent masked tokens. We used DNABERT-2’s BPE tokenizer, trained on 2.75 billion nucleotide bases from the human nuclear genome and 32.49 billion bases from 135 species across various kingdoms. Also, we trained custom BPE tokenizers with varying vocabulary sizes using our DNA barcode dataset and evaluated across various training scenarios on the genus-level 1-NN probing task. Based on results in Table 4, we selected a loss weight of 1.0 since it yielded the best results. Similarly, our setup followed from Table 4, as we maintain a 50% substitution token ratio ($r_{s}$). All the results for the various configurations can be found in Table 5.
Model size. We report genus-level 1-NN probing accuracy on the unseen data for DNABERT-2 BPE and BarcodeBERT BPE tokenizers across three model sizes. Models using DNABERT-2’s BPE showed no significant accuracy changes across different sizes. However, for BarcodeBERT BPE, we see that increasing the model size reduces the accuracy across different vocabulary sizes, possibly due to overfitting.
Vocabulary size. We trained new BPE tokenizers with varying vocabulary sizes ($v$) to evaluate their impact on BarcodeBERT BPE performance. Unlike $k$-mer tokenizers, BPE sequence lengths vary with input composition and require padding ([UNK]) or truncation to a maximum length before tokenization. Since smaller vocabularies ($v$) produce longer sequences (see Table 11), we set maximum lengths of 128 for $v=4096,1024$ and 256 for $v=256,128$. Reducing the vocabulary size from 4,096 to 128 consistently decreases accuracy across all model sizes.
Comparing $k$-mer with BPE. We find that $k$-mer tokenizers outperform BPE tokenizers in all model configurations. We hypothesize that this is likely due to three reasons. First, DNA barcode sequences are too short to benefit from the BPE compression. Second, BPE is sensitive to minor variations such as single-character substitutions, which is unsuitable for DNA datasets with single-nucleotide mutations in arbitrary positions. In other words, BPE tokenization varies greatly for similar sequences with small Hamming distances, while $k$-mer tokenization remains consistent, as a single-nucleotide substitution affects only one token (see Table 11). Third, although BPE handles small sequence alignment shifts better than $k$-mer tokenizers, this can be mitigated in $k$-mers with data augmentation using random offsets during pretraining (see Table 12).
@@ 5 Discussion
Our results demonstrate that pretraining masked language models on DNA barcode data, as exemplified by BarcodeBERT, is highly effective for arthropod species identification. Our pretrained model performs well on various downstream tasks common in biodiversity analyses, such as taxonomic classification, clustering, and similarity searches. BarcodeBERT excels in these tasks because it efficiently uses hardware acceleration, enabling it to scale effectively for large datasets, while being faster (55x) than alignment-based approaches in species level classification. By systematically evaluating tokenization and masking strategies, we also provide actionable insights for the pretraining of DNA-specific foundation models.
Despite its strengths, BarcodeBERT has some limitations. For instance, its training data may have taxonomic and geographical biases. In particular, the model is trained exclusively on invertebrate species from Canada, potentially limiting its applicability in global studies. In this context, the BOLD dataset, now comprising more than 16 million DNA barcodes from a wide geographical distribution (Ratnasingham & Hebert, 2007), represents a wealth of untapped data that could address these biases. Future work should incorporate more diverse datasets to develop robust, globally scalable self-supervised models for taxonomic classification. The methodology and findings we present should be broadly applicable to barcode regions for other kingdoms, such as the ITS region for fungi (Schoch et al., 2012), however the method is yet to be validated beyond the COI barcode region used for Animalia.
Lastly, while genomic sequences longer than barcodes could offer deeper insights for specialized phylogenetic analyses, the quadratic time complexity of transformer models limits their applicability to such sequences. Future work should include more parameter-efficient architectures such as Structured State Space Models, which scale sub-quadratically with sequence length (Gu et al., 2022). These architectures may be pretrained on specific barcode datasets to provide a more efficient alternative for applications involving longer sequences.
@@ 6 Conclusions
Overall, BarcodeBERT’s performance demonstrates how transformer-based architectures can be successfully customized to overcome the challenges of genomic biodiversity data, for effective DNA barcode identification and classification. Lastly, not being limited to a specific dataset or barcode region, our model is highly amenable to future applications, to global datasets or barcode regions in other kingdoms of life.
@@ Appendix A Dataset Description
After pre-processing (see main text), the final dataset contained 965,289 samples. The dataset was divided into three main partitions: Pretrain, Seen, and Unseen. The Seen partition was subdivided into supervised training, validation, and testing subsets. Each partition serves distinct experimental purposes, with the Unseen partition designed to mimic real-world scenarios where models encounter sequences from previously unobserved species.
Table 6 summarizes the composition of each partition, detailing the number of unique records across taxonomic categories, from phylum to species. It also includes the Barcode Index Number (BIN), a species-level proxy formed by clustering similar DNA barcode sequences into a molecular taxonomic unit. The Pretrain partition contains 893,744 sequences from 15 phyla, while the Seen and Unseen partitions are more specific, containing sequences primarily from phylum Arthropoda. The Unseen partition contains 4,278 barcodes from 1,826 species absent from the training and validation subsets.
@@
Pretrain 15 49 176 1188 6679 14794 893744 62489
Test 1 1 13 161 500 1653 13460 2125
Train 1 1 13 161 500 1653 47086 2430
Validation 1 1 13 161 500 1653 6721 1954
Unseen 1 1 13 161 500 1826 4278 1885
@@
The partitions provide a structured dataset that supports the evaluation of models in both closed- and open-world settings. It enables a robust assessment of models’ ability to classify known taxa and generalize to new species.
@@ B.1 Variation in Length of Tokenized Sequences
Table 11 shows the minimum, mean, and maximum lengths of tokenized sequences in DNA barcode pretraining data. Generally, BPE tokenizers with smaller vocabulary sizes tend to produce longer tokenized sequences, with the only exception being the DNABERT-2 BPE. According to Table 11, although the tokenizer trained on DNABERT-2 uses a vocabulary size of 4096, it generates relatively long tokenized sequences. This is because the dataset used to train DNABERT-2 differs from the DNA barcode dataset used to create other tokenizers.
Table 11: Comparison of tokenized sequence lengths in the DNA barcode pretraining data across different BPE tokenizers with varying vocabulary sizes.
Tokenizer vocabulary sizes ($v$) Min Mean Max
DNABERT-2 4096 48 126.95 398
BarcodeBERT 4096 20 77.17 388
1024 30 104.61 407
256 54 146.59 429
128 65 175.44 451
@@ B.2 Sensitivity of Tokenizers to DNA Sequence Variations
There are two main ways to measure similarity between sequences. One is through the Hamming distance, which counts the number of mismatches between two strings, i.e., the number of positions where the strings do not share the same nucleotide. The other is the edit distance, also called Levenshtein distance, which is defined as the minimum number of string operations (substitutions, insertions, deletions) required to transform one string into the other. In comparative genomics, the edit distance is widely used to estimate evolutionary relationships, as it corresponds to a better model of true sequence similarity. However, the Hamming distance can be calculated in linear time, even in token space with larger vocabularies. Figure 7 illustrates how the Hamming distance varies with a single insertion operation versus a single substitution operation in both sequence space and token space for a pair of sequences. We include the two tokenization strategies considered in our study: $k$-mer and DNABERT-2 BPE tokenization. In both scenarios, the edit distance is equal to one, but just the BPE tokenizer can model this behaviour using the Hamming distance as a proxy.
@@ Appendix C Masking Strategy
In this extra experiment, we adapted our masking strategy to mimic BERT’s original methodology (Devlin et al., 2019). BERT addresses the difference in token distribution between pretraining and fine-tuning by employing a masking strategy whereby 80% of substitution tokens are replaced by [MASK] tokens, 10% are replaced with random tokens, and 10% remain unchanged. This approach ensures more robust embeddings during testing, where masked tokens are absent (Devlin et al., 2019).
@@ D.2 Baseline Models
For evaluation, we utilized the respective Pretrained models from Huggingface ModelHub, specifically:
D.3 Pretraining

### v3

@@ 4.2.2 Zero-shot image classification using DNA embeddings
Table 2: Evaluation of DNA barcode models in a Bayesian zero-shot learning task on the INSECT dataset. The pretraining and fine-tuning data source is indicated by the respective DNA type, and ‘–’ signifies the absence of training for that type. We also indicate the most specific taxon subset. For the baseline CNN encoder, we report the original paper result (left) and our reproduction (right). Numbers in boldface indicate the best result across each task, and underlined second best.
@@ 4.3 Ablation Studies
In our main results, we demonstrated the utility of self-supervised pretraining BarcodeBERT to enable our model to generalize to open-world tasks. In this section, we study the impact of the different components involved in pretraining. We use the terminology context tokens for tokens that are left unchanged to provide context to the model during pretraining and the terminology substitution tokens for tokens that will be changed for the masked language modelling task. We consider different strategies to calculate the loss of each group separately. The loss associated with predicting contextual tokens is referred to as the “context component” of the loss [^1] Although predictions in the foundation model literature are typically restricted to substitution tokens, we extend this to include context token predictions, maintaining the terminology to explore their utility in input reconstruction similar to autoencoders., while the loss related to predicting substitution tokens is the “substitution component”. By assigning different weights to these two loss components, we sought to observe how these adjustments would affect both training and evaluation. In particular, we define $w_{s}$ as the penalty weight given to the substitution component of the loss. Note that $1\!-\!w_{s}$ is always the weight of the context component of the loss.
@@ 4.3.2 Weight of the substitution component of the loss
Table 3: Classification accuracy over the different substitution token ratios $r_{s}$, while keeping constant all of the model architecture (4-4), the value of $k\!=\!4$ during tokenization and the penalty weight for the substitution component of the loss ($w_{s}\!=\!1$). Numbers in boldface indicate the highest accuracies and the italic value shows the $\rightarrow$ selected $\leftarrow$ shows the selected optimal parameter.
@@
Fine-tuned Linear probe 1-NN probe
@@
Table 4: Genus-level accuracy for 1-NN probing of unseen species with varying penalty weight assigned to the substitution component of the loss ($w_{s}$). The model architecture remains fixed (4-4), with substitution token ratio $r_{s}$ = 50%. Two tokenizers were tested: a $k$-mer tokenizer with $k$-mer sizes of 2, 4, 6, and 8, and a BPE tokenizer used in DNABERT-2 with a fixed vocabulary size of 4096. Note that the weight of the context component of the loss always equals $1-w_{s}$. The number in boldface indicates the overall best accuracy, underlined the best per tokenizer, and italic the $\rightarrow$ selected $\leftarrow$ optimal parameter.
@@ 4.3.3 Tokenization strategies
We evaluated the BPE tokenizer, which generates variable-length tokens based on character co-occurrence frequencies (Sennrich et al., 2016). BPE was designed for subword tokenization to overcome fixed vocabulary (Sennrich et al., 2016) by compressing sequences for efficient representation (Gallé, 2019). Unlike overlapping $k$-mers, BPE has an advantage for masked DNA language models (Zhou et al., 2023) as it avoids information leaks from adjacent masked tokens. We used DNABERT-2’s BPE tokenizer, trained on 2.75 billion nucleotide bases from the human nuclear genome and 32.49 billion bases from 135 species across various kingdoms. We also trained custom BPE tokenizers with varying vocabulary sizes on our DNA barcode dataset and evaluated across various training scenarios on the genus-level 1-NN probing task. Based on results in Table 12, we selected a loss weight of 1.0 for best performance. Our setup used a 50% substitution token ratio ($r_{s}$). Full results across configurations are shown in Table 5.
Model size. We report genus-level 1-NN probing accuracy on the unseen data for DNABERT-2 BPE and BarcodeBERT BPE tokenizers across three model sizes. Models using DNABERT-2’s BPE showed no significant accuracy changes across different sizes. However, for BarcodeBERT BPE, we see that increasing the model size reduces the accuracy across different vocabulary sizes, possibly due to overfitting. For additional experiments, see Supplementary Materials D.3.
Vocabulary size. We trained new BPE tokenizers with varying vocabulary sizes ($v$) to evaluate their impact on BarcodeBERT BPE performance. Unlike $k$-mer tokenizers, BPE sequence lengths vary with input composition and require padding or truncation to a maximum length before tokenization. Since smaller vocabularies ($v$) produce longer sequences (see Supplement B.1), we set maximum lengths of 128 for $v\!\in\!\{4096,1024\}$ and 256 for $v\!\in\!\{256,128\}$. Reducing the vocabulary size from 4096 to 128 consistently decreases accuracy across all model sizes.
Comparing $k$-mer with BPE. We find that $k$-mer tokenizers outperform BPE tokenizers in all model configurations. We hypothesize that this is likely due to three reasons. First, DNA barcode sequences are too short to benefit from the BPE compression. Second, BPE is sensitive to minor variations such as single-character substitutions, which is unsuitable for DNA datasets with single-nucleotide mutations in arbitrary positions. In other words, BPE tokenization varies greatly for similar sequences with small Hamming distances, while $k$-mer tokenization remains consistent, as a single-nucleotide substitution affects only one token (see Supplementary Figure S9). Third, although BPE handles small sequence alignment shifts better than $k$-mer tokenizers, this can be mitigated in $k$-mers with data augmentation using random offsets during pretraining (see Supplementary Table S6).
@@ 5 Discussion
Our results demonstrate that pretraining masked language models on DNA barcode data, as exemplified by BarcodeBERT, is highly effective for arthropod species identification. Despite only  35% of the pretraining data being labelled at the species rank, the model learns latent features that purely supervised approaches cannot capture. BarcodeBERT performs well on key biodiversity tasks, such as taxonomic classification, clustering, and similarity searches, benefiting from efficient hardware acceleration that enables scaling for large datasets and achieving $\times5$ faster species-level classification than alignment-based methods. By systematically evaluating tokenization and masking strategies, we also provide actionable insights for the pretraining of DNA-specific foundation models.
Despite its strengths, BarcodeBERT has some limitations. Its training data may have taxonomic and geographical biases as the model is trained exclusively on invertebrate species from Canada, potentially limiting its applicability in global studies. The BOLD dataset, comprising more than 16 million DNA barcodes from a wide geographical distribution (Ratnasingham & Hebert, 2007), represents untapped data that could address such biases. Future work should incorporate more diverse datasets to develop robust, globally scalable models for taxonomic classification. Our methodology and findings should be broadly applicable to barcode regions for other kingdoms, such as the ITS region for fungi (Schoch et al., 2012), however, the method requires validation beyond the COI barcode region used for Animalia.
Lastly, while longer genomic sequences could offer deeper insights for specialized phylogenetic analyses, the quadratic time complexity of transformer models limits their applicability to such sequences. Future work should include more computationally efficient architectures such as structured state space models, which scale sub-quadratically with sequence length (Gu et al., 2022).
@@ 6 Conclusions
Overall, BarcodeBERT’s performance demonstrates how transformer-based architectures can be successfully customized to overcome the challenges of genomic biodiversity data for effective DNA barcode identification and classification. Lastly, not being limited to a specific dataset or barcode region, our model is highly amenable to future applications, to global datasets or barcode regions in other kingdoms of life.
@@ Appendix A Dataset Description
After pre-processing (see subsubsection 3.1.1), the final dataset contained 965,289 samples. The dataset was divided into three main partitions: Pretrain, Seen, and Unseen. The Seen partition was subdivided into supervised training, validation, and testing subsets. Each partition serves distinct experimental purposes, with the Unseen partition designed to mimic real-world scenarios where models encounter sequences from previously unobserved species. Our preprocessed version of the dataset is available to download at https://huggingface.co/datasets/bioscan-ml/CanadianInvertebrates-ML.
Table 6 summarizes the composition of each partition, detailing the number of unique records across taxonomic categories, from phylum to species. It also includes the number of unique Barcode Index Numbers (BINs); each BIN serves as a species-level proxy, created by clustering similar DNA barcode sequences into a molecular taxonomic unit. The Pretrain partition contains 893,744 sequences from 15 phyla, while the Seen and Unseen partitions are more specific, containing sequences primarily from phylum Arthropoda. The Unseen partition contains 4,278 barcodes from 1,826 species absent from the training and validation subsets.
@@
Pretrain 15 49 176 1,188 6,679 14,794 893,744 62,489
Test 1 1 13 161 500 1,653 13,460 2,125
Train 1 1 13 161 500 1,653 47,086 2,430
Validation 1 1 13 161 500 1,653 6,721 1,954
Unseen 1 1 13 161 500 1,826 4,278 1,885
@@
Lastly, we analyzed the variability of COI DNA barcodes across insect species in the test partition using the normalized edit distance and cosine distance on the learned embeddings. Only species with more than two sequences are considered, and for each species $\mathcal{S}$, we computed both intra- and inter-species distances. Specifically, for each sequence, we calculated:
  • The distance to all other sequences within the same species (intra-species), and
  • The distance to all sequences from different species (inter-species),
while ensuring distances were not double-counted. Finally, we estimated each distribution via Gaussian kernel density estimation (KDE) and computed overlap by integrating the minimum of the two KDE curves.
Figure 7: Distribution of intra- and inter-species distances based on normalized edit distance computed on the test partition. Blue represents intra-species distances, orange represents inter-species distances, and the area of overlap is shaded in green. Distances were calculated using Levenshtein alignment, normalized using min-max normalization. Species with at least two sequences were included in the analysis.
The edit distance (Figure 7) provides a clear separation between intra- and inter-species distributions, with minimal overlap (0.15). These distributions quantitatively capture the natural variability in COI sequences and illustrate the potential for learning taxonomic structure with deep models. This intuition is confirmed by the results obtained after estimating the distribution of the cosine distances of the learned embeddings (Figure 8). Indeed, self-supervised pretraining enables BarcodeBERT to encode biologically relevant distinctions more clearly than raw alignment metrics with an overlap of 0.02 between both distributions. While these distances reflect biological structure and are useful in our $k$-NN classifier and zero-shot classification (ZSC) settings, the separation between intra- and inter-species groups should be further investigated, specially constrained to other taxonomic ranks.
Figure 8: Distribution of intra- and inter-species distances based on cosine distance computed on GAP embeddings generated by the model. Blue represents intra-species distances, orange represents inter-species distances, and the area of overlap is shaded in green. While cosine distances on learned embeddings capture biological structure, the separation between intra- and inter-species groups is less distinct than with alignment-based metrics.
@@ B.1 Variation in Length of Tokenized Sequences
Figure 9 shows the distribution of tokenized sequence lengths in DNA barcode pretraining data. Generally, BPE tokenizers with smaller vocabulary sizes tend to produce longer tokenized sequences, with the only exception being the DNABERT-2 BPE. According to Figure 9, although the tokenizer trained on DNABERT-2 uses a vocabulary size of 4096, it generates relatively long tokenized sequences. This is because the dataset used to train DNABERT-2 differs from the DNA barcode dataset used to create other tokenizers.
Figure 9: Distribution of tokenized sequence lengths across different BPE tokenizers, shown as histograms for the DNA barcode pretraining data. The distributions correspond to different BPE tokenizers, including DNABERT-2 and BarcodeBERT with vocabulary sizes ($v$) of 4096, 1024, 256, and 128. Smaller vocabulary sizes result in longer tokenized sequences, as reflected by the rightward shift in the distribution. Summary statistics (minimum, mean, and maximum lengths) are provided in the table inset.
@@ B.2 Sensitivity of Tokenizers to DNA Sequence Variations
There are two main ways to measure similarity between sequences. One is through the Hamming distance, which counts the number of mismatches between two strings, i.e., the number of positions where the strings do not share the same nucleotide. The other is the edit distance, also called the Levenshtein distance, which is defined as the minimum number of string operations (substitutions, insertions, deletions) required to transform one string into the other. In comparative genomics, the edit distance is widely used to estimate evolutionary relationships, as it corresponds to a better model of true sequence similarity. However, the Hamming distance can be calculated in linear time, even in token space with larger vocabularies. Figure 10 illustrates how the Hamming distance varies with a single insertion operation versus a single substitution operation in both sequence space and token space for a pair of sequences. We include the two tokenization strategies considered in our study: $k$-mer and DNABERT-2 BPE tokenization. In both scenarios, the edit distance is equal to one, but just the BPE tokenizer can model this behaviour using the Hamming distance as a proxy.
@@ Appendix C Masking Strategy
In this supplementary experiment, we adapted our masking strategy to mimic BERT’s original methodology (Devlin et al., 2019). BERT addresses the difference in token distribution between pretraining and fine-tuning by employing a masking strategy whereby 80% of substitution tokens are replaced by [MASK] tokens, 10% are replaced with random tokens, and 10% remain unchanged. This approach ensures more robust embeddings during testing, where masked tokens are absent (Devlin et al., 2019).
@@ D.2 Baseline Models
For evaluation, we utilized the respective pretrained models from Hugging Face’s ModelHub, specifically:
D.3 Model architecture configuration
To identify the optimal model configuration, we conducted a series of ablation experiments evaluating the impact of architectural choices, tokenization strategies, and the substitution component of the loss ($w_{s}$). We considered $k$-mer tokenization with four distinct $k$ values (2, 4, 6, 8), as well as the BPE tokenizer. Three transformer configurations were evaluated: (i) 4 layers with 4 attention heads, (ii) 6 layers with 6 heads, and (iii) 12 layers with 12 heads. Additionally, we varied the penalty weight assigned to the substitution component of the loss ($w_{s}$) function, exploring values from 0.2 to 1.0.
Table 12 reports genus-level accuracy for 1-NN probing of unseen species across these settings. The best performance was obtained with the 4-4-4 configuration (4-mers, 4 heads, 4 layers), which we adopt as the default model throughout our evaluations.
Table 12: Genus-level classification accuracy for 1-NN probing of unseen species using different $k$-mer sizes, transformer configurations, and substitution loss weights ($w_{s}$). Best results for each configuration are indicated in boldface.
Genus-level acc (%)
of unseen species with $1$-NN probe
4 layers, 4 heads 6 layers, 6 heads 12 layers, 12 heads
Loss weight ($w_{s}$) $k\!=\!2$ $k\!=\!4$ $k\!=\!6$ $k\!=\!8$ BPE $k\!=\!2$ $k\!=\!4$ $k\!=\!6$ $k\!=\!8$ BPE $k\!=\!2$ $k\!=\!4$ $k\!=\!6$ $k\!=\!8$ BPE
0.2 64.18 76.06 75.15 71.15 70.57 61.59 74.61 70.87 67.74 67.15 48.92 63.72 57.12 56.40 62.34
0.5 66.47 74.98 76.62 71.22 70.34 65.38 73.37 70.87 69.70 67.57 46.23 67.11 61.24 60.05 62.20
0.8 68.84 76.71 74.66 73.33 69.40 68.37 74.02 72.20 69.72 68.23 60.50 67.50 66.87 61.05 67.09
1.0 76.92 78.47 75.74 75.62 69.85 67.71 73.91 74.38 75.33 70.45 73.98 68.27 68.67 73.79 68.16
D.4 Pretraining